OpenStudy (g0thw1ng):

Write the point-slope form of the line satisfying the conditions. Then use the point-slope form of the equation to write the slope-intercept form of the equation in function notation. Slope = - 2/3, passing through (0, 5) Select one: a. f(x) = 2/3x - 5 b. f(x) = -2/3 x - 5 c. f(x) = - 2/3x + 5 d. f(x) = -2/3 x - 15/2

OpenStudy (sloppycanada):

The equation you're looking for is y=mx+b If 'm' is considered the slope, and 2/3 = slope for the problem,m, then the value of m is?

OpenStudy (g0thw1ng):

-2/3 is the m then what do I do with the (0,5)

OpenStudy (welshfella):

point=slope form is y - 5 = -(2/3)(x - 0) which you can transform into slope intercept form

OpenStudy (welshfella):

the general form of slope point is y - y1 = m(x - x1) where (x1,y1) = a point on the line In this case x1 = 0 and y1 = 5